Find the number of ordered pairs (x;y) of positive integers satisfying 1/x + 1/y = 1/(2021 ^ 17)


If x ≠ y then pairs (x;y) and (y;x) are considered to be different.


1
Expert's answer
2021-11-15T17:45:31-0500

"\frac{x+y}{xy}=\frac{1}{2021^{17}}"


"xy-2021^{17}(x+y)=0"


Adding "(2021^{17})^2" both sides


"xy-2022^{17}(x+y)+(2021^{17})^2=(2021^{17})^2"


"(x-2021^{17})(y-2021^{17})=(2021^{17})^2"


Let "x-2021^{17}=A" and "y-2021^{17}=B"


"\therefore\>AB=(2021^{17})^2"


"(2021^{17})^2=43^{34}\u00d747^{34}"


Number of factors "=" "(34+1)(34+1)"

"=1225"


In one case "A=B"


"\therefore" Number of ordered pairs "=1224"